OpenStudy (anonymous):

Can someone please help me translate stuff on a graph? I will medal and fan!

OpenStudy (anonymous):

I have to translate some equations and write a story about it, and I can't quite figure it out… f(x)=-|4x+6|+4 is the first one and it needs to be moved 4 to the left

OpenStudy (anonymous):

-abs(4x+10)+4

OpenStudy (anonymous):

Thank you sooooo much! I don't know why I couldn't figure that out /).(\ gosh I'm stupid today

OpenStudy (anonymous):

Any chance you could help me with a few more?

OpenStudy (anonymous):

\[y≥\frac{ 1 }{ 2 }x^2 -2\]

OpenStudy (anonymous):

\[y \ge \frac{ 1 }{ 2 }x ^{2} -2\]

OpenStudy (anonymous):

that one needs to be moved 4 to the left also

OpenStudy (anonymous):

@alex122is

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

my mind is a piece of floppy jelly right now

OpenStudy (anonymous):

just put +4 like the other one..... 1/2(x+4)^2-2

OpenStudy (anonymous):

Thanks again…

OpenStudy (anonymous):

you are free to go now XD

OpenStudy (anonymous):

bye